### v2

# Movable Signals with Dual-Polarized Fixed Intelligent Surfaces: Beyond Diagonal Reflection Matrices

Abstract
This paper investigates wireless systems aided by dual-polarized intelligent surfaces. We compare reconfigurable intelligent surface (RIS), which adjust their reflection matrices, with movable signals operating with fixed intelligent surface (FIS), which adjust the signal frequency while the surface properties remain fixed. For both RIS and FIS, we consider surfaces with a diagonal reflection matrix, named diagonal RIS/FIS, and surfaces with a reflection matrix not limited to being diagonal, named beyond-diagonal RIS/FIS. Movable signals with FIS always outperform RIS, achieving at least a fourfold gain. When transmitter and receiver polarizations differ, beyond-diagonal FIS further enhances performance.
I Introduction
Reconfigurable intelligent surfaces (RISs) have been thoroughly investigated as a promising means to enable smart radio environments (SREs) by shaping and controlling the wireless channel [1, 2]. A RIS is a surface composed of reflecting elements with adjustable electromagnetic (EM) properties, thereby steering incident signals toward desired directions. Although RIS can offer interesting performance gains, its practical deployment remains challenging. Dynamic reconfiguration of the RIS requires dedicated control circuitry with non-negligible power consumption. An effective RIS reconfiguration relies on accurate channel knowledge, demanding additional training overhead. Moreover, since a RIS alters all incident EM signals, the resulting unpredictable interference from different operators raises concerns.
To circumvent these limitations, an alternative approach has been introduced: movable signals operating with fixed intelligent surface (FIS) [3]. Instead of reconfiguring the EM properties of the surface, this approach reconfigures the carrier frequency of the transmitted signal, while using a surface of uniformly spaced reflecting elements with fixed EM properties. Prior work has demonstrated that, in single-input single-output (SISO) systems, movable signals with FIS can enhance the channel strength by up to four times compared with RIS [3]. Although this gain relies on a suitable spectrum availability (ideally, the available carrier frequencies should be distributed over a wide range), it remarkably does not require any reconfigurable hardware or high channel estimation overhead.
Previous work investigated FIS with a diagonal reflection matrix, which we denote as diagonal FIS (D-FIS), mirroring the assumption adopted for conventional RIS, also known as diagonal RIS (D-RIS) [3, 4]. Nevertheless, recent developments in beyond-diagonal RIS (BD-RIS) highlight the benefits of artificially coupling the RIS elements to each other, which leads to more flexible reflection capabilities. Extending this concept to FIS and exploring the performance benefits of beyond-diagonal FIS (BD-FIS), i.e., FISs with fixed reflection matrices not limited to be diagonal, remains unexplored.
In this work, we investigate a practical version of FIS where the antenna elements are dual-polarized, and show that BD-FIS can be particularly beneficial in this case. Dual-polarization has been introduced in multiple-input multiple-output (MIMO) communications to allow more antennas within a limited space, and increase diversity [5], and it is reasonable to expect that it will also be considered for intelligent surfaces. Focusing on SISO systems aided by dual-polarized intelligent surfaces, we analyze and compare the performance of four designs: D-RIS, BD-RIS, D-FIS, and BD-FIS, where FISs are used jointly with movable signals.
Our contributions are the following. First, we characterize the received power achievable with these four designs in the two cases where the transmitter and receiver have the same and opposite polarization. In this analysis, we consider the effect of the specular reflection due to the structural scattering of the surface, which is often neglected in the RIS literature, but plays an important role [6, 7, 8]. Second, we obtain in closed form the performance gain of BD-RIS, D-FIS, and BD-FIS over D-RIS. Movable signals with FIS always outperform RIS, achieving at least a four times higher received power. Furthermore, when the transmitter and receiver have opposite polarization, BD-FIS yields even larger performance improvements, reaching infinite gains as the inverse of the cross-polar discrimination (XPD) approaches zero.
II Dual-Polarized RIS/FIS-Aided Channel Model and Problem Formulation
Consider a SISO communication system, as illustrated in Fig. 1. To enhance the channel strength, a reflecting surface made of $N$ dual-polarized elements is deployed, which can be either a RIS or a FIS, depending on whether its reflecting properties are reconfigurable or fixed. Letting the transmitted signal be $x\in\mathbb{C}$, the received signal writes as $y=hx+n$, where $h\in\mathbb{C}$ is the wireless channel and $n\in\mathbb{C}$ is the noise.
II-A Channel Model
Following previous work [1, 3], the channel $h$ is given by $h=\mathbf{h}_{R}\boldsymbol{\Theta}\mathbf{h}_{T}-\mathbf{h}_{R}\mathbf{h}_{T}$, where $\boldsymbol{\Theta}\in\mathbb{C}^{N\times N}$ is the reflection matrix of the $N$-element surface, $\mathbf{h}_{R}\in\mathbb{C}^{1\times N}$ is the channel between the surface and receiver, and $\mathbf{h}_{T}\in\mathbb{C}^{N\times 1}$ is the channel between the transmitter and surface. In $h$, the term $\mathbf{h}_{R}\boldsymbol{\Theta}\mathbf{h}_{T}$ accounts for the signal reflected by the surface and controlled by $\boldsymbol{\Theta}$, while the term $\mathbf{h}_{R}\mathbf{h}_{T}$ accounts for the signal specularly reflected by the surface (an effect of the structural scattering which is independent of $\boldsymbol{\Theta}$ [6, 7, 8]). The direct link between the transmitter and receiver is considered to be very weak and hence omitted.
Figure 1: Dual-polarized RIS/FIS-aided system where the transmitter and receiver have opposite polarization.
The channels $\mathbf{h}_{R}$ and $\mathbf{h}_{T}$ are modeled by assuming that the surface is a dual-polarized uniform linear array (ULA) in line-of-sight (LoS) with both the transmitter and receiver. The $N$ elements of the surface are numbered such that the first $N/2$ have vertical polarization and the last $N/2$ have horizontal polarization, as shown in Fig. 1. The effect of the element polarization is captured in $\mathbf{h}_{R}$ and $\mathbf{h}_{T}$ by modeling them as
$\mathbf{h}_{R}=\mathbf{p}_{R}\otimes\mathbf{g}_{R},\;\mathbf{h}_{T}=\mathbf{p}_{T}\otimes\mathbf{g}_{T},$ (1)
where $\otimes$ is the Kronecker product, $\mathbf{p}_{R}\in\mathbb{R}^{1\times 2}$ and $\mathbf{p}_{T}\in\mathbb{R}^{2\times 1}$ capture the power imbalance between the two polarizations, and $\mathbf{g}_{R}\in\mathbb{C}^{1\times N/2}$ and $\mathbf{g}_{T}\in\mathbb{C}^{N/2\times 1}$ are the uni-polarized LoS fading channels [5]. The vectors $\mathbf{p}_{R}$ and $\mathbf{p}_{T}$ depend on the polarizations of the transmitter and receiver antennas. For a vertically or horizontally polarized receiver, $\mathbf{p}_{R}=[1,\sqrt{\chi}]$ or $\mathbf{p}_{R}=[\sqrt{\chi},1]$, respectively, where $0\leq\chi\leq 1$ is the inverse of the XPD. Likewise, for a vertically or horizontally polarized transmitter, $\mathbf{p}_{T}=[1,\sqrt{\chi}]^{T}$ or $\mathbf{p}_{T}=[\sqrt{\chi},1]^{T}$, respectively.
The LoS fading channels $\mathbf{g}_{R}$ and $\mathbf{g}_{T}$ are expressed as follows. We place the $N$ elements of the surface on the $x$-axis, centered in $x=0$, so that the $n$th element has $x$ coordinate $x_{n}=(n-(N+1)/2)d_{A}$, with $d_{A}$ being the spacing between the antenna elements. Denoting as $d_{R}$ the distance between the center of the surface and the receiver, and as $\theta_{R}\in[-\pi/2,\pi/2]$ the angle of the receiver with respect to the surface normal, the distance between the $n$th element and the receiver is $d_{R,n}=d_{R}-x_{n}\sin(\theta_{R})$ assuming far-field. Hence, the entries of $\mathbf{g}_{R}$, given by $\mathbf{g}_{R}]_{n}=e^{-j2\pi d_{R,n}/\lambda}$, are
$\left[\mathbf{g}_{R}\right]_{n}=e^{-j\frac{2\pi}{\lambda}\left[d_{R}-\left(n-\frac{N+1}{2}\right)d_{A}\sin(\theta_{R})\right]},$ (2)
as a function of the wavelength $\lambda=c/f$, where $c$ the speed of light, and $f$ the frequency. Likewise, denoting as $d_{T}$ the distance between the center of the surface and the transmitter, and as $\theta_{T}\in[-\pi/2,\pi/2]$ the angle of the transmitter, the entries of $\mathbf{g}_{T}$ are written as
$\left[\mathbf{g}_{T}\right]_{n}=e^{-j\frac{2\pi}{\lambda}\left[d_{T}-\left(n-\frac{N+1}{2}\right)d_{A}\sin(\theta_{T})\right]}.$ (3)
II-B Problem Formulation
In this communication system, our metric is the received power, written as
$P_{R}=P_{T}\left|\mathbf{h}_{R}\boldsymbol{\Theta}\mathbf{h}_{T}-\mathbf{h}_{R}\mathbf{h}_{T}\right|^{2},$ (4)
where $P_{T}=\mathbb{E}[|x|^{2}]$ is the transmitted power, assumed $P_{T}=1$ for simplicity. We compare two strategies for maximizing $P_{R}$.
II-B1 RIS
With a RIS, the received power is maximized by reconfiguring $\boldsymbol{\Theta}$ depending on the channels $\mathbf{h}_{R}$ and $\mathbf{h}_{T}$, which are assumed to be known [1]. Formally, the problem is
$\underset{\boldsymbol{\Theta}}{\mathsf{\mathrm{max}}}\;\;P_{R}\;\;\mathsf{\mathrm{s.t.}}\;\;\mathbf{h}_{R}\text{ and }\mathbf{h}_{T}\text{ are fixed},$ (5)
where $\boldsymbol{\Theta}$ is reconfigurable subject to specific constraints. For a D-RIS, we have $\boldsymbol{\Theta}=\text{diag}\left(e^{j\theta_{1}},\ldots,e^{j\theta_{N}}\right)$, where $\theta_{n}\in[0,2\pi)$ are the phase shifts that can be reconfigured. For a BD-RIS, $\boldsymbol{\Theta}$ can be any unitary matrix, i.e., $\boldsymbol{\Theta}^{H}\boldsymbol{\Theta}=\mathbf{I}$, if only the lossless constraint is considered.
II-B2 Movable Signals with FIS
With movable signals, the received power is maximized by reconfiguring the frequency $f$ (or the wavelength $\lambda$), while the reflection matrix of the surface $\boldsymbol{\Theta}$ is fixed to a suitable predefined value. Such a surface with a fixed reflection matrix is referred to as FIS [3]. The optimization problem for movable signals is
[…]
where $\boldsymbol{\Theta}$ is pre-optimized offline and fixed during deployment. For a D-FIS, $\boldsymbol{\Theta}$ is set subject to $\boldsymbol{\Theta}=\text{diag}\left(e^{j\theta_{1}},\ldots,e^{j\theta_{N}}\right)$, while for a BD-FIS, it is subject to $\boldsymbol{\Theta}^{H}\boldsymbol{\Theta}=\mathbf{I}$.
In the two sections that follow, we compare the received power achievable through RIS and movable signals with FIS, by separately considering the two cases where the transmitter and receiver have the same and opposite polarization.
III Same Polarization at the Tx and Rx
Assume the transmitter and receiver to have the same polarization, namely vertical, with no loss of generality. Following (1), this yields $\mathbf{h}_{R}=[\mathbf{g}_{R},\sqrt{\chi}\mathbf{g}_{R}]$ and $\mathbf{h}_{T}=[\mathbf{g}_{T}^{T},\sqrt{\chi}\mathbf{g}_{T}^{T}]^{T}$.
III-1 Diagonal RIS
With a D-RIS, it is well known that the received power is maximized by setting $\boldsymbol{\Theta}$ as $\boldsymbol{\Theta}=\text{diag}(e^{j\theta_{1}},\ldots,e^{j\theta_{N}})$, with phase shifts $\theta_{n}=-\arg([\mathbf{h}_{R}]_{n}[\mathbf{h}_{T}]_{n})+\arg(-\mathbf{h}_{R}\mathbf{h}_{T})$, for $n=1,\dots,N$. This global optimal solution gives
$P_{R}^{\mathrm{DRIS}}=\left(\sum_{n=1}^{N}\left|\left[\mathbf{h}_{R}\right]_{n}\left[\mathbf{h}_{T}\right]_{n}\right|+\left|\mathbf{h}_{R}\mathbf{h}_{T}\right|\right)^{2}.$ (7)
Since $\mathbf{h}_{R}=[\mathbf{g}_{R},\sqrt{\chi}\mathbf{g}_{R}]$ and $\mathbf{h}_{T}=[\mathbf{g}_{T}^{T},\sqrt{\chi}\mathbf{g}_{T}^{T}]^{T}$, we have $|[\mathbf{h}_{R}]_{n}[\mathbf{h}_{T}]_{n}|=1$, for $n=1,\ldots,N/2$, $|[\mathbf{h}_{R}]_{n}[\mathbf{h}_{T}]_{n}|=\chi$, for $n=N/2+1,\ldots,N$, and $\mathbf{h}_{R}\mathbf{h}_{T}=(1+\chi)\mathbf{g}_{R}\mathbf{g}_{T}$, giving
$P_{R}^{\mathrm{DRIS}}$ $=\frac{\left(1+\chi\right)^{2}}{4}\left(N+2\left|\mathbf{g}_{R}\mathbf{g}_{T}\right|\right)^{2}.$ (8)
Taking the expectation of (8) over $\mathbf{g}_{R}$ and $\mathbf{g}_{T}$, whose entries are assumed to have independent phases uniformly distributed in $0,2\pi)$, the average received power $\bar{P}_{R}^{\mathrm{DRIS}}=\mathbb{E}[P_{R}^{\mathrm{DRIS}}]$ is
$\bar{P}_{R}^{\mathrm{DRIS}}=\frac{\left(1+\chi\right)^{2}}{4}\left(N^{2}+4N\mathbb{E}\left[\left|\mathbf{g}_{R}\mathbf{g}_{T}\right|\right]+4\mathbb{E}[\left|\mathbf{g}_{R}\mathbf{g}_{T}\right|^{2}]\right).$ (9)
Furthermore, since $\mathbf{g}_{R}\mathbf{g}_{T}$ is the sum of $N/2$ independent and identically distributed (i.i.d.) random variables with mean 0 and variance 1, it can be approximated as a complex Gaussian with mean 0 and variance $N/2$, i.e., $\mathbf{g}_{R}\mathbf{g}_{T}\sim\mathcal{CN}(0,N/2)$, by the Central Limit Theorem. Hence, by exploiting $\mathbb{E}[|\mathbf{g}_{R}\mathbf{g}_{T}|]=\sqrt{\pi N/8}$ and $\mathbb{E}[|\mathbf{g}_{R}\mathbf{g}_{T}|^{2}]=N/2$, we obtain
$\bar{P}_{R}^{\mathrm{DRIS}}=\frac{\left(1+\chi\right)^{2}}{4}\left(N^{2}+\sqrt{2\pi N}N+2N\right),$ (10)
showing that the received power with a D-RIS when transmitter and receiver have the same polarization scales with $\bar{P}_{R}^{\mathrm{DRIS}}\approx(1+\chi)^{2}N^{2}/4$.
III-2 Beyond-Diagonal RIS
With a BD-RIS, the achievable received power is given by
$P_{R}^{\mathrm{BDRIS}}=\left(\left\|\mathbf{h}_{R}\right\|\left\|\mathbf{h}_{T}\right\|+\left|\mathbf{h}_{R}\mathbf{h}_{T}\right|\right)^{2},$ (11)
which can be achieved by reconfiguring $\boldsymbol{\Theta}$ with the global optimal solution derived in [9]. Since $\mathbf{h}_{R}=[\mathbf{g}_{R},\sqrt{\chi}\mathbf{g}_{R}]$ and $\mathbf{h}_{T}=[\mathbf{g}_{T}^{T},\sqrt{\chi}\mathbf{g}_{T}^{T}]^{T}$, we have $\|\mathbf{h}_{R}\|^{2}=\|\mathbf{h}_{T}\|^{2}=(1+\chi)N/2$ and $\mathbf{h}_{R}\mathbf{h}_{T}=(1+\chi)\mathbf{g}_{R}\mathbf{g}_{T}$, giving that $P_{R}^{\mathrm{BDRIS}}$ is equal to the received power with D-RIS in (8). Therefore, the average received power achievable with a BD-RIS $\bar{P}_{R}^{\mathrm{BDRIS}}=\mathbb{E}[P_{R}^{\mathrm{BDRIS}}]$ is the same as the average received power with a D-RIS given by (10), scaling with $\bar{P}_{R}^{\mathrm{BDRIS}}\approx(1+\chi)^{2}N^{2}/4$. The fact that D-RIS and BD-RIS achieve the same performance in dual-polarized systems when the transmitter and receiver have the same polarization was also derived in [10], although the specular reflection was neglected in [10].
III-3 Diagonal FIS
With movable signals and D-FIS, the reflection matrix of the FIS is pre-optimized offline subject to $\boldsymbol{\Theta}=\text{diag}\left(e^{j\theta_{1}},\ldots,e^{j\theta_{N}}\right)$, while the signal frequency $f$ is dynamically reconfigured. We show how to fix $\boldsymbol{\Theta}$ and optimize $f$ by first deriving an upper bound on the received power, and then giving solutions for $\boldsymbol{\Theta}$ and $f$ which achieve it. The received power with a D-FIS is upper bounded by
$P_{R}$ $\leq\left(\sum_{n=1}^{N}\left|\left[\mathbf{h}_{R}\right]_{n}\left[\mathbf{h}_{T}\right]_{n}\right|+\left|\mathbf{h}_{R}\mathbf{h}_{T}\right|\right)^{2}$ (12)
$=\left(1+\chi\right)^{2}\left(\frac{N}{2}+\left|\mathbf{g}_{R}\mathbf{g}_{T}\right|\right)^{2}\leq\left(1+\chi\right)^{2}N^{2},$ (13)
where (12) is given by the triangle inequality, and (13) follows from $|[\mathbf{h}_{R}]_{n}[\mathbf{h}_{T}]_{n}|=1$, for $n=1,\ldots,N/2$, $|[\mathbf{h}_{R}]_{n}[\mathbf{h}_{T}]_{n}|=\chi$, for $n=N/2+1,\ldots,N$, $\mathbf{h}_{R}\mathbf{h}_{T}=(1+\chi)\mathbf{g}_{R}\mathbf{g}_{T}$, and $|\mathbf{g}_{R}\mathbf{g}_{T}|\leq N/2$. To achieve the upper bound $P_{R}=(1+\chi)^{2}N^{2}$, the D-FIS matrix is fixed to $\boldsymbol{\Theta}^{\star}=-\mathbf{I}_{N}$, such that the term $\mathbf{h}_{R}\boldsymbol{\Theta}\mathbf{h}_{T}$ is aligned with the specular reflection $\mathbf{h}_{R}\mathbf{h}_{T}$. With $\boldsymbol{\Theta}=-\mathbf{I}_{N}$, the received power is
$P_{R}=\left|-\mathbf{h}_{R}\mathbf{h}_{T}-\mathbf{h}_{R}\mathbf{h}_{T}\right|^{2}=4\left(1+\chi\right)^{2}\left|\mathbf{g}_{R}\mathbf{g}_{T}\right|^{2},$ (14)
since $\mathbf{h}_{R}\mathbf{h}_{T}=(1+\chi)\mathbf{g}_{R}\mathbf{g}_{T}$, and therefore the signal frequency needs to be reconfigured to maximize $|\mathbf{g}_{R}\mathbf{g}_{T}|$, i.e., to ensure that $|\mathbf{g}_{R}\mathbf{g}_{T}|=N/2$. Interestingly, this is achieved if the signal wavelength is reconfigured to $\lambda^{\star}=d_{A}|\sin(\theta_{R})+\sin(\theta_{T})|$, depending on the angles $\theta_{R}$ and $\theta_{T}$, with corresponding frequency
$f^{\star}=\frac{c}{d_{A}\left|\sin(\theta_{R})+\sin(\theta_{T})\right|}.$ (15)
The optimality of this value can be verified by substituting $\lambda^{\star}$ into the expressions of $\mathbf{g}_{R}$ and $\mathbf{g}_{T}$ in (2) and (3) and verifying that it leads to $|\mathbf{g}_{R}\mathbf{g}_{T}|=N/2$. Therefore, by reconfiguring the signal frequency as in (15) a D-FIS with reflection matrix $\boldsymbol{\Theta}=-\mathbf{I}_{N}$ can achieve a received power $P_{R}^{\mathrm{DFIS}}=(1+\chi)^{2}N^{2}$. Note that this value is four times higher than the received power achieved by a D-RIS $\bar{P}_{R}^{\mathrm{DRIS}}\approx(1+\chi)^{2}N^{2}/4$, and can be obtained without reconfiguring the EM properties of the surface. Movable signals can achieve higher performance than RIS since they control both terms $\mathbf{h}_{R}\boldsymbol{\Theta}\mathbf{h}_{T}$ and $\mathbf{h}_{R}\mathbf{h}_{T}$ in the received power expression, while RIS only controls $\mathbf{h}_{R}\boldsymbol{\Theta}\mathbf{h}_{T}$.
III-4 Beyond-Diagonal FIS
With movable signals and BD-FIS, the reflection matrix of the FIS is pre-optimized offline subject to $\boldsymbol{\Theta}^{H}\boldsymbol{\Theta}=\mathbf{I}$, while the signal frequency $f$ is dynamically reconfigured to maximize the received power. In this case, the received power is upper bounded by
$P_{R}$ $\leq\left(\left\|\mathbf{h}_{R}\right\|\left\|\mathbf{h}_{T}\right\|+\left|\mathbf{h}_{R}\mathbf{h}_{T}\right|\right)^{2}$ (16)
$=\left(1+\chi\right)^{2}\left(\frac{N}{2}+\left|\mathbf{g}_{R}\mathbf{g}_{T}\right|\right)^{2}\leq\left(1+\chi\right)^{2}N^{2},$ (17)
where (16) is due to the triangle inequality and the sub-multiplicity of the $\ell_{2}$-norm, and (17) holds because $\|\mathbf{h}_{R}\|^{2}=\|\mathbf{h}_{T}\|^{2}=(1+\chi)N/2$, $\mathbf{h}_{R}\mathbf{h}_{T}=(1+\chi)\mathbf{g}_{R}\mathbf{g}_{T}$, and $|\mathbf{g}_{R}\mathbf{g}_{T}|\leq N/2$. Therefore, the maximum received power achievable by a BD-FIS is the same as the received power of a D-FIS given by (13). This implies that $\boldsymbol{\Theta}=-\mathbf{I}$ is global optimal also for a BD-FIS, and the signal frequency can still be reconfigured as in (15) to achieve a received power of $P_{R}^{\mathrm{BDFIS}}\approx(1+\chi)^{2}N^{2}$.
We have studied a dual-polarized RIS/FIS-aided system where the transmitter and receiver have the same polarization, considering D-RIS, BD-RIS, and movable signals with D-FIS and BD-FIS. The scaling laws of the received power in these four cases are summarized in Tab. I, from which we make two observations. First, BD-RIS and BD-FIS do not improve the received power over D-RIS and D-FIS, respectively. Second, movable signals with FIS can obtain four times higher received power than fixed-frequency signals with RIS.
IV Opposite Polarization at the Tx and Rx
Assume now the transmitter and receiver to have opposite polarization, namely vertical and horizontal, respectively, with no loss of generality. Following (1), this yields $\mathbf{h}_{R}=[\sqrt{\chi}\mathbf{g}_{R},\mathbf{g}_{R}]$ and $\mathbf{h}_{T}=[\mathbf{g}_{T}^{T},\sqrt{\chi}\mathbf{g}_{T}^{T}]^{T}$.
IV-1 Diagonal RIS
With a D-RIS, the maximum received power is given by (7). Since now the channels are $\mathbf{h}_{R}=[\sqrt{\chi}\mathbf{g}_{R},\mathbf{g}_{R}]$ and $\mathbf{h}_{T}=[\mathbf{g}_{T}^{T},\sqrt{\chi}\mathbf{g}_{T}^{T}]^{T}$, we have $|[\mathbf{h}_{R}]_{n}[\mathbf{h}_{T}]_{n}|=\sqrt{\chi}$, for $n=1,\ldots,N$, and $\mathbf{h}_{R}\mathbf{h}_{T}=2\sqrt{\chi}\mathbf{g}_{R}\mathbf{g}_{T}$, allowing us to simplify (7) as
$P_{R}^{\mathrm{DRIS}}$ $=\chi\left(N+2\left|\mathbf{g}_{R}\mathbf{g}_{T}\right|\right)^{2}.$ (18)
Taking the expectation of (18) over all possible channels $\mathbf{g}_{R}$ and $\mathbf{g}_{T}$, the average received power $\bar{P}_{R}^{\mathrm{DRIS}}=\mathbb{E}[P_{R}^{\mathrm{DRIS}}]$ is
$\bar{P}_{R}^{\mathrm{DRIS}}=\chi\left(N^{2}+4N\mathbb{E}\left[\left|\mathbf{g}_{R}\mathbf{g}_{T}\right|\right]+4\mathbb{E}[\left|\mathbf{g}_{R}\mathbf{g}_{T}\right|^{2}]\right),$ (19)
where we can make the approximation $\mathbf{g}_{R}\mathbf{g}_{T}\sim\mathcal{CN}(0,N/2)$, following the Central Limit Theorem. Therefore, by using $\mathbb{E}[|\mathbf{g}_{R}\mathbf{g}_{T}|]=\sqrt{\pi N/8}$ and $\mathbb{E}[|\mathbf{g}_{R}\mathbf{g}_{T}|^{2}]=N/2$, we get
$\bar{P}_{R}^{\mathrm{DRIS}}=\chi\left(N^{2}+\sqrt{2\pi N}N+2N\right),$ (20)
giving the scaling law of the received power in the presence of a D-RIS as $\bar{P}_{R}^{\mathrm{DRIS}}\approx\chi N^{2}$.
TABLE I: Scaling laws of the received power in
dual-polarized RIS/FIS-aided systems.
Same polariz. at Tx and Rx Opposite polariz. at Tx and Rx
$\begin{aligned} \bar{P}_{R}^{\mathrm{DRIS}}&=\left(1+\chi\right)^{2}N^{2}/4\\ \bar{P}_{R}^{\mathrm{BDRIS}}&=\left(1+\chi\right)^{2}N^{2}/4\\ P_{R}^{\mathrm{DFIS}}&=\left(1+\chi\right)^{2}N^{2}\\ P_{R}^{\mathrm{BDFIS}}&=\left(1+\chi\right)^{2}N^{2}\end{aligned}$ $\begin{aligned} \bar{P}_{R}^{\mathrm{DRIS}}&=\chi N^{2}\\ \bar{P}_{R}^{\mathrm{BDRIS}}&=\left(1+\chi\right)^{2}N^{2}/4\\ P_{R}^{\mathrm{DFIS}}&=4\chi N^{2}\\ P_{R}^{\mathrm{BDFIS}}&=\left(1+\chi+2\sqrt{\chi}\right)^{2}N^{2}/4\end{aligned}$
IV-2 Beyond-Diagonal RIS
In the presence of a BD-RIS, the maximum received power is given by (11). Since $\mathbf{h}_{R}=[\sqrt{\chi}\mathbf{g}_{R},\mathbf{g}_{R}]$ and $\mathbf{h}_{T}=[\mathbf{g}_{T}^{T},\sqrt{\chi}\mathbf{g}_{T}^{T}]^{T}$ give $\|\mathbf{h}_{R}\|^{2}=\|\mathbf{h}_{T}\|^{2}=(1+\chi)N/2$ and $\mathbf{h}_{R}\mathbf{h}_{T}=2\sqrt{\chi}\mathbf{g}_{R}\mathbf{g}_{T}$, (11) becomes
$P_{R}^{\mathrm{BDRIS}}=\left(\left(1+\chi\right)\frac{N}{2}+2\sqrt{\chi}\left|\mathbf{g}_{R}\mathbf{g}_{T}\right|\right)^{2}.$ (21)
Taking the expectation of (21) over all possible channels $\mathbf{g}_{R}$ and $\mathbf{g}_{T}$, the average received power $\bar{P}_{R}^{\mathrm{BDRIS}}=\mathbb{E}[P_{R}^{\mathrm{BDRIS}}]$ is
$\bar{P}_{R}^{\mathrm{BDRIS}}=\frac{\left(1+\chi\right)^{2}}{4}N^{2}+2\left(1+\chi\right)\sqrt{\chi}N\mathbb{E}\left[\left|\mathbf{g}_{R}\mathbf{g}_{T}\right|\right]\\ +4\chi\mathbb{E}[\left|\mathbf{g}_{R}\mathbf{g}_{T}\right|^{2}].$ (22)
Furthermore, by approximating $\mathbf{g}_{R}\mathbf{g}_{T}$ as $\mathbf{g}_{R}\mathbf{g}_{T}\sim\mathcal{CN}(0,N/2)$ following the Central Limit Theorem, we can apply $\mathbb{E}[|\mathbf{g}_{R}\mathbf{g}_{T}|]=\sqrt{\pi N/8}$ and $\mathbb{E}[|\mathbf{g}_{R}\mathbf{g}_{T}|^{2}]=N/2$ to obtain
$\bar{P}_{R}^{\mathrm{BDRIS}}=\frac{\left(1+\chi\right)^{2}}{4}N^{2}+\left(1+\chi\right)\sqrt{\frac{\pi}{2}\chi N}N+2\chi N,$ (23)
showing that the received power with a BD-RIS scales with $\bar{P}_{R}^{\mathrm{BDRIS}}\approx(1+\chi)^{2}N^{2}/4$, always higher than $\bar{P}_{R}^{\mathrm{DRIS}}\approx\chi N^{2}$ when $\chi\neq 1$.
IV-3 Diagonal FIS
When considering movable signals and D-FIS, the received power is upper bounded by
$P_{R}$ $\leq\left(\sum_{n=1}^{N}\left|\left[\mathbf{h}_{R}\right]_{n}\left[\mathbf{h}_{T}\right]_{n}\right|+\left|\mathbf{h}_{R}\mathbf{h}_{T}\right|\right)^{2}$ (24)
$=\left(\sqrt{\chi}N+2\sqrt{\chi}|\mathbf{g}_{R}\mathbf{g}_{T}|\right)^{2}\leq 4\chi N^{2},$ (25)
where (24) is given by the triangle inequality, and (25) follows from $|[\mathbf{h}_{R}]_{n}[\mathbf{h}_{T}]_{n}|=1$, for $n=1,\ldots,N/2$, $|[\mathbf{h}_{R}]_{n}[\mathbf{h}_{T}]_{n}|=\chi$, for $n=N/2+1,\ldots,N$, $\mathbf{h}_{R}\mathbf{h}_{T}=2\sqrt{\chi}\mathbf{g}_{R}\mathbf{g}_{T}$, and $|\mathbf{g}_{R}\mathbf{g}_{T}|\leq N/2$. As discussed in Section III-3, the upper bound $P_{R}=4\chi N^{2}$ can be achieved by fixing the D-FIS matrix to $\boldsymbol{\Theta}^{\star}=-\mathbf{I}_{N}$. With $\boldsymbol{\Theta}=-\mathbf{I}_{N}$, the received power is
$P_{R}=\left|-\mathbf{h}_{R}\mathbf{h}_{T}-\mathbf{h}_{R}\mathbf{h}_{T}\right|^{2}=16\chi\left|\mathbf{g}_{R}\mathbf{g}_{T}\right|^{2},$ (26)
since $\mathbf{h}_{R}\mathbf{h}_{T}=2\sqrt{\chi}\mathbf{g}_{R}\mathbf{g}_{T}$, and therefore the signal frequency needs to be optimized to maximize $|\mathbf{g}_{R}\mathbf{g}_{T}|$. As discussed in Section III-3, $|\mathbf{g}_{R}\mathbf{g}_{T}|$ is maximized when the signal frequency is set as in (15), depending on the angles $\theta_{R}$ and $\theta_{T}$. Therefore, by using a signal frequency as in (15) a D-FIS with reflection matrix $\boldsymbol{\Theta}=-\mathbf{I}_{N}$ can achieve a received power $P_{R}^{\mathrm{DFIS}}=4\chi N^{2}$. Note that this value is four times higher than the received power achieved by a D-RIS $\bar{P}_{R}^{\mathrm{DRIS}}\approx\chi N^{2}$, while requiring no reconfiguration of EM properties of the surface.
Figure 2: (a) A phase shifter with phase $\pi$ as an example of a 22-port network with scattering matrix $0,-1]^{T},[-1,0]^{T}]$, and (b) implementation of the $n$th group of the BD-FIS in Section IV-4.
IV-4 Beyond-Diagonal FIS
With movable signals and BD-FIS, the received power is upper bounded by
$P_{R}$ $\leq\left(\left\|\mathbf{h}_{R}\right\|\left\|\mathbf{h}_{T}\right\|+\left|\mathbf{h}_{R}\mathbf{h}_{T}\right|\right)^{2}$ (27)
$=\left(\left(1+\chi\right)\frac{N}{2}+2\sqrt{\chi}\left|\mathbf{g}_{R}\mathbf{g}_{T}\right|\right)^{2}\leq\frac{\left(1+\chi+2\sqrt{\chi}\right)^{2}}{4}N^{2},$ (28)
where (27) is due to the triangle inequality and the sub-multiplicity of the $\ell_{2}$-norm, and (28) holds since $\|\mathbf{h}_{R}\|^{2}=\|\mathbf{h}_{T}\|^{2}=(1+\chi)N/2$, $\mathbf{h}_{R}\mathbf{h}_{T}=2\sqrt{\chi}\mathbf{g}_{R}\mathbf{g}_{T}$, and $|\mathbf{g}_{R}\mathbf{g}_{T}|\leq N/2$. We now show how to fix the BD-FIS matrix and optimize the signal frequency to achieve this upper bound. The BD-FIS matrix is fixed to
$\boldsymbol{\Theta}=\begin{bmatrix}\mathbf{0}_{N/2}&-\mathbf{I}_{N/2}\\ -\mathbf{I}_{N/2}&\mathbf{0}_{N/2}\end{bmatrix}$, (29)
which means we are considering a group-connected architecture with $N/2$ groups, each including 2 elements [11] (note that (29) is a block diagonal matrix with $2\times 2$ blocks being $0,-1]^{T},[-1,0]^{T}]$ if its columns are permuted). The groups are formed by two co-located elements having opposite polarization and all have a $2\times 2$ scattering matrix given by $0,-1]^{T},[-1,0]^{T}]$. Since an example of $2$-port network having scattering matrix $0,-1]^{T},[-1,0]^{T}]$ is a phase shifter with phase $\pi$ (Fig. 2(a)), such a BD-FIS can be implemented by connecting each pair of co-located elements via a phase shifter with phase $\pi$ (Fig. 2(b)). In this way, the signal incident on the vertically polarized elements is reflected from the horizontally polarized ones, and vice versa, and the BD-FIS is a polarization converter. With (29), the received power is
$P_{R}$ $=\left|\mathbf{h}_{R}\begin{bmatrix}\mathbf{0}_{N/2}&-\mathbf{I}_{N/2}\\ -\mathbf{I}_{N/2}&\mathbf{0}_{N/2}\end{bmatrix}\mathbf{h}_{T}-\mathbf{h}_{R}\mathbf{h}_{T}\right|^{2}$ (30)
$=\left(1+\chi+2\sqrt{\chi}\right)^{2}\left|\mathbf{g}_{R}\mathbf{g}_{T}\right|^{2},$ (31)
where we used $\mathbf{h}_{R}=[\sqrt{\chi}\mathbf{g}_{R},\mathbf{g}_{R}]$ and $\mathbf{h}_{T}=[\mathbf{g}_{T}^{T},\sqrt{\chi}\mathbf{g}_{T}^{T}]^{T}$. The signal frequency needs now to be reconfigured to maximize $|\mathbf{g}_{R}\mathbf{g}_{T}|$, i.e., to ensure that $|\mathbf{g}_{R}\mathbf{g}_{T}|=N/2$, which can be achieved by setting it as in (15). Therefore, by reconfiguring the signal frequency as in (15) a BD-FIS with reflection matrix given by (29) achieves a received power $P_{R}^{\mathrm{BDFIS}}=(1+\chi+2\sqrt{\chi})^{2}N^{2}/4$. Remarkably, this value is higher than the received power achieved by a D-FIS $P_{R}^{\mathrm{DFIS}}=4\chi N^{2}$ when $\chi\neq 1$, and always higher than the received power of BD-RIS.
We have studied a dual-polarized RIS/FIS-aided system where the transmitter and receiver have opposite polarization. The scaling laws of the received power achieved when considering D-RIS, BD-RIS, and movable signals with D-FIS and BD-FIS are summarized in Tab. I. We observe that BD-RIS and BD-FIS improve the received power over D-RIS and D-FIS when $\chi\neq 1$, respectively, and that movable signals with FIS always outperform fixed-frequency signals with RIS.
Figure 3: Gain of BD-RIS, D-FIS, and BD-RIS over D-RIS as a function of $\chi$, when the transmitter and receiver have opposite polarization.
V Gain over Diagonal RIS
In this section, we quantify the gain of BD-RIS, D-FIS, and BD-FIS over D-RIS, defined as $G^{\mathrm{X}}=\lim_{N\to+\infty}\bar{P}_{R}^{\mathrm{X}}/\bar{P}_{R}^{\mathrm{DRIS}}$, where $\mathrm{X}\in\{\mathrm{BDRIS},\;\mathrm{DFIS},\;\mathrm{BDFIS}\}$. When the transmitter and receiver have the same polarization, these gains are given by $G^{\mathrm{BDRIS}}=1$, $G^{\mathrm{DFIS}}=4$, and $G^{\mathrm{BDFIS}}=4$, which be shown by using the scaling laws in (10), (13), and (17). Besides, when the transmitter and receiver have opposite polarization, these gains are $G^{\mathrm{BDRIS}}=(1+\chi)^{2}/(4\chi)$, $G^{\mathrm{DFIS}}=4$, and $G^{\mathrm{BDFIS}}=(1+\chi+2\sqrt{\chi})^{2}/(4\chi)$, which are obtained by using the scaling laws in (20), (23), (25), and (28).
Focusing on systems where the transmitter and receiver have opposite polarization, we report the gains $G^{\mathrm{BDRIS}}$, $G^{\mathrm{DFIS}}$, and $G^{\mathrm{BDFIS}}$ as a function of $\chi$ in Fig. 3. We make the following three observations. First, the gain of BD-RIS over D-RIS tends to infinity when $\chi=0$, decreases with $\chi$, and is $1$ when $\chi=1$. BD-RIS is therefore particularly beneficial for small values of $\chi$, as also explained in [10]. Second, the gain of D-FIS over D-RIS is always $4$ and does not vary with $\chi$, as for uni-polarized systems [3]. Third, the gain of BD-FIS over D-RIS tends to infinity when $\chi=0$, decreases with $\chi$, and reaches $4$ when $\chi=1$. Therefore, BD-FIS with movable signals is an appealing strategy for achieving substantially higher performance than RIS in dual-polarized systems. Since its reconfigurability lies in the signal frequency, BD-FIS has a fixed reflection matrix, which is easy to implement and does not require control hardware.
VI Conclusion
We have investigated the performance limits of SISO systems aided by dual-polarized intelligent surfaces. Specifically, we have compared two strategies: RIS, where the reflection properties of the surface are reconfigured, and movable signals with FIS, where the signal frequency is reconfigured while the surface has fixed reflection properties. We show that movable signals with FIS always outperform RIS, by achieving at least a four times higher received power. When the transmitter and receiver have opposite polarization, a FIS with a reflection matrix not limited to be diagonal, namely a BD-FIS, remarkably further increases the performance. Future research could explore BD-FIS architectures that are optimal for any polarization of the transmitter and receiver, and the optimization of the frequency of movable signals under realistic spectrum constraints and channel models, e.g., in the near-field.